OpenStudy (anonymous):

find the value of x. 1/5- x = 3

OpenStudy (luigi0210):

First what you do is move everything to one side, so move that 1/5 over

OpenStudy (anonymous):

-x=3 *5/1 is that correct?

OpenStudy (luigi0210):

Is the equation: \[\frac{1}{5}-x=3\] OR \[\frac{1}{5}x=3\]

OpenStudy (anonymous):

and you multiply both sides by 5/1 to get the x alone right?

OpenStudy (luigi0210):

It depends on which which equation it is

OpenStudy (anonymous):

its is the first one you posted

OpenStudy (luigi0210):

Alright. Since it's +/- we can't multiply. We have to do the opposite operation

OpenStudy (anonymous):

and by that you mean?

OpenStudy (luigi0210):

Mean if the numer is positive we subtract, of it's negative we add it. Like this one: